OpenStudy (anonymous):

HELP PLEASE !!! Find and simplify the difference quotient f(x+h)−f(x) / h for the given function. f(x)=1/5x

OpenStudy (tkhunny):

First, let's fix your notation. You do not mean what you have typed: f(x+h)−f(x) / h You mean this: [f(x+h)−f(x)] / h Rememeber your Order of Operations and that typing "in-line" is a little different from text book printing. Second, you'll have to specify a function. There is no general procedure for all functions.

OpenStudy (anonymous):

i forgot 1 thing :( sorry

OpenStudy (anonymous):

f(x)=1/5x

OpenStudy (tkhunny):

Okay. What does f(x+h) mean, knowing f(x) = 1/5x? By the way, do you mean \[\frac{ 1 }{ 5x }\] or \[\frac{ 1 }{ 5 } \cdot x\] ? You wrote the second one.

OpenStudy (anonymous):

\[\frac{ 1 }{ 5x }\]

OpenStudy (tkhunny):

Okay. What does f(x+h) mean, knowing f(x) = 1/(5x)?

OpenStudy (anonymous):

that's what they asked me , i don't understand

OpenStudy (anonymous):

:)